OpenStudy (anonymous):

Suppose the diameter of a circle is 7. What is its circumference?

OpenStudy (anonymous):

@haleyelizabeth2017

OpenStudy (haleyelizabeth2017):

isn't circumference \[2 \pi r^2\]?

OpenStudy (haleyelizabeth2017):

so if so.....then it is \(2 \pi (3.5)^2\)

OpenStudy (haleyelizabeth2017):

then you do \((3.5)^2\) first which is 12.25 then multiply that by 2. which equals \(24.50 \pi\) or 77 (rounded up)

myininaya (myininaya):

Circumference of a circle= \[C=2 \pi r \text{ or use } C=\pi d\] where r is radius and d is diameter

OpenStudy (haleyelizabeth2017):

oh....whoopsies I haven't used it in a while so I forgot :(